# Robustness-Inspired Defense Against Backdoor Attacks on Graph Neural Networks

### v1

@@ G.4 Implementation Details
Appendix H Results on Clean Graph
@@
In this section, we conduct experiments to show the clean accuracy of our RIGBD when trained on a clean graph. Specifically, we follow the experimental setup described in Sec. 5.2 and remove the backdoor triggers from the poisoned graph. Then, we train a GCN and RIGBD on this clean graph. The results are shown in Tab. 6. From the table, we observe that our RIGBD achieves comparable clean accuracy to GCN. This is because our poisoned node selection strategy, defined in Eq. (5), stops when nodes continuously show different labels compared to the node with the highest prediction variance. Since nodes with high confidence are not always from the same class in a clean graph, we rarely select nodes as poisoned nodes. This highlights the adaptability of our RIGBD, even when we do not know whether the graph is clean or poisoned.
Appendix I Additional Results of Defense Performance
@@
Appendix J Additional Results of the Ability to Detect Poisoned Nodes
@@
In this section, we provide additional results of the ability to detect poisoned nodes on Cora and PubMed dataset. We present the recall and precision of RIGBD in identifying poisoned nodes. We use the same setting as in Sec. 5.2. The results are shown in Table 3. We also present the corresponding ASR and ACC to illustrate the correlation between detection ability and defense performance. Clean accuracy (Clean ACC), obtained from a model trained on the clean graph, is provided as a reference to evaluate the model’s performance on clean nodes. From the table, we observe the following: (i) Our RIGBD demonstrates consistently high precision, always over $90\%$, in detecting poisoned nodes across three attack methods in all datasets, with detection recall always exceeding $80\%$. This indicates that our strategy of random edge dropping with graph convolution operation defined in Eq. (3) typically leads to higher prediction variance for poisoned nodes. (ii) Although the recall for some attack methods and datasets is less than $90\%$, we still achieve an ASR close to $0\%$ while maintaining ACC. This demonstrates the stability of our strategy to train a robust node classifier using Eq. (6). Even when some of the poisoned nodes are not identified, it still efficiently helps the model unlearn the triggers.
Appendix K Additional Results of Hyperparameter Analysis
@@
In this section, we provide additional results on how different drop ratios ($\beta$) and different numbers of iterations of random edge dropping ($K$) impact the ACC of RIGBD. ASR, Recall, and Precision are also provided as references. Specifically, we vary the values of $K$ as $\left\{2,5,25,50,100\right\}$, and the values of $\beta$ as $\left\{0.1,0.3,0.5,0.7,0.9\right\}$. The attack method used is DPGBA. The other settings are the same as the evaluation protocol in Sec. 5.1. The results are shown in Fig. 9. From the figure, we make the following observations: (i) Our RIGBD is stable in terms of ASR, ACC, and Precision. Notably, when the drop ratio $\beta=0.1$ and we only conduct two iterations of random edge dropping, the recall of our poison node detection is around 17.4%. However, we still achieve robust defense performance with ASR close to 0%. This demonstrates that our robust training strategy is effective even with a lower recall. As our framework identifies the most efficient triggers that lead to large prediction variance and focuses on minimizing their impact. By doing so, the influence of less efficient triggers is inherently counteracted. In contrast, when we simply remove those 17.4% triggers from the dataset and train a model on the remaining dataset, the ASR is around 80%. (ii) As $K$ increases, the recall also increases, empirically verifying our Theorem 1 that in expectation, the clean nodes will have stable node representations and thus stable predictions when using the graph convolution operation designed in Eq. (3). As $\beta$ increases from 0.1 to 0.7, the recall also increases without a decrease in precision, demonstrating the stability of our method. When $\beta=0.9$, only about 10% of edges remain in each iteration. Though the recall decreases slightly, we still achieve consistently high precision. This further indicates the robustness of our model against various chosen hyperparameters.
Appendix L Additional Results of Ablation Studies
@@
Appendix M Additional Results of Defense Performance and Ability to Detect Poison Nodes with Different Numbers of Triggers
@@
In this section, we conduct experiments to demonstrate how varying numbers of backdoor triggers impact the performance of RIGBD in terms of backdoor defense and poisoned node detection. Specifically, we set the number of triggers to $\left\{10,20,40,80,160\right\}$ for Cora and $\left\{20,80,160,240,320\right\}$ for PubMed. The attack method used is DPGBA. The drop ratio $\beta=0.5$, and the number of iterations for random edge dropping are set as $K=10$ for Cora and $K=20$ for PubMed. The results are shown in Table 9. From the table, we observe: (i) Though there are fluctuations in the recall and precision in detecting poisoned nodes, the precision always exceeds $85\%$. This indicates that our method of random edge dropping consistently leads to higher prediction variance, helping us precisely identify poisoned nodes regardless of the number of triggers. (ii) With the different number of triggers, our method demonstrate consistently superior performance in defense, with ASR close to $0\%$ and ACC comparable to Clean ACC. Notably, when the number of triggers is $240$ for PubMed, the detection recall is around $66\%$, but we still achieve good results in terms of ASR and ACC. This further indicates the stability of our method in training a robust node classifier using Eq. (6), even if some poisoned nodes are not identified. Additional results on the visualization of prediction variance for poisoned nodes and clean nodes with different numbers of triggers are provided in Fig. 11 and Fig. 12.

### v2

@@ H.5 Implementation details
Appendix I Results on Clean Graph
@@
In this section, we conduct experiments to show the clean accuracy of our RIGBD when trained on a clean graph. Specifically, we follow the experimental setup described in Sec. 5.2 and remove the backdoor triggers from the poisoned graph. Then, we train a GCN and RIGBD on this clean graph. The results are shown in Tab. 7. From the table, we observe that our RIGBD achieves comparable clean accuracy to GCN. This is because our poisoned node selection strategy, defined in Eq. 5, stops when nodes continuously show different labels compared to the node with the highest prediction variance. Since nodes with high confidence are not always from the same class in a clean graph, we rarely select nodes as poisoned nodes. This highlights the adaptability of our RIGBD, even when we do not know whether the graph is clean or poisoned.
Appendix J Additional Results of Defense Performance
@@
Appendix K Additional Results of Defense Performance on Heterophilic Graphs
Dataset Methods Clean ACC ASR Recall Precision
Wisconsin GCN 48 100 - -
RIGBD 48 0 90 95
Cornell GCN 50.0 93.3 - -
RIGBD 53.5 3.3 70 100
Texas GCN 44.4 92.9 - -
RIGBD 42.9 3.7 80 70
Table 9: Performance comparison of GCN and RIGBD across different datasets.
In this section, we conduct additional experiments to evaluate the performance of RIGBD on heterophilic graphs. The attack method used is DPGBA Zhang et al. (2024), with 20 poisoned nodes during training, and other settings follow those described in Section 5.1. The results are presented in Table 9. It is evident that RIGBD achieves consistently strong performance in defending against backdoor attacks while maintaining clean accuracy. The recall and precision of poisoned node detection further demonstrate that our method remains robust when applied to heterophilic graphs.
Appendix L Additional Results of the Ability to Detect Poisoned Nodes
@@
Physics GTA 95.8 0.32 95.10 86.2 91.8
UGBA 95.8 0.12 95.71 96.0 98.4
DPGBA 95.8 0.21 95.97 96.8 99.2
Flickr GTA 44.4 0.00 44.21 69.5 98.1
UGBA 44.4 0.00 42.74 98.6 97.2
DPGBA 44.4 0.00 43.80 97.9 100.0
@@
In this section, we provide additional results of the ability to detect poisoned nodes on Cora and PubMed dataset. We present the recall and precision of RIGBD in identifying poisoned nodes. We use the same setting as in Sec. 5.2. The results are shown in Table 10. We also present the corresponding ASR and ACC to illustrate the correlation between detection ability and defense performance. Clean accuracy (Clean ACC), obtained from a model trained on the clean graph, is provided as a reference to evaluate the model’s performance on clean nodes. From the table, we observe the following: (i) Our RIGBD demonstrates consistently high precision, always over $90\%$, in detecting poisoned nodes across three attack methods in all datasets, with detection recall always exceeding $80\%$. This indicates that our strategy of random edge dropping with graph convolution operation defined in Eq. 3 typically leads to higher prediction variance for poisoned nodes. (ii) Although the recall for some attack methods and datasets is less than $90\%$, we still achieve an ASR close to $0\%$ while maintaining ACC. This demonstrates the stability of our strategy to train a robust node classifier using Eq. 6. Even when some of the poisoned nodes are not identified, it still efficiently helps the model unlearn the triggers.
Appendix M Additional Results of Hyperparameter Analysis
@@
In this section, we provide additional results on how different drop ratios ($\beta$) and different numbers of iterations of random edge dropping ($K$) impact the ACC of RIGBD. ASR, Recall, and Precision are also provided as references. Specifically, we vary the values of $K$ as $\left\{2,5,25,50,100\right\}$, and the values of $\beta$ as $\left\{0.1,0.3,0.5,0.7,0.9\right\}$. The attack method used is DPGBA. The other settings are the same as the evaluation protocol in Sec. 5.1. The results are shown in Fig. 9. From the figure, we make the following observations: (i) Our RIGBD is stable in terms of ASR, ACC, and Precision. Notably, when the drop ratio $\beta=0.1$ and we only conduct two iterations of random edge dropping, the recall of our poison node detection is around 17.4%. However, we still achieve robust defense performance with ASR close to 0%. This demonstrates that our robust training strategy is effective even with a lower recall. As our framework identifies the most efficient triggers that lead to large prediction variance and focuses on minimizing their impact. By doing so, the influence of less efficient triggers is inherently counteracted. In contrast, when we simply remove those 17.4% triggers from the dataset and train a model on the remaining dataset, the ASR is around 80%. (ii) As $K$ increases, the recall also increases, empirically verifying our Theorem 1 that in expectation, the clean nodes will have stable node representations and thus stable predictions when using the graph convolution operation designed in Eq. 3. As $\beta$ increases from 0.1 to 0.7, the recall also increases without a decrease in precision, demonstrating the stability of our method. When $\beta=0.9$, only about 10% of edges remain in each iteration. Though the recall decreases slightly, we still achieve consistently high precision. This further indicates the robustness of our model against various chosen hyperparameters.
Appendix N Additional Results of Ablation Studies
@@
Appendix O Additional Results of Defense Performance and Ability to Detect Poison Nodes with Different Numbers of Triggers
@@
In this section, we conduct experiments to demonstrate how varying numbers of backdoor triggers impact the performance of RIGBD in terms of backdoor defense and poisoned node detection. Specifically, we set the number of triggers to $\left\{10,20,40,80,160\right\}$ for Cora and $\left\{20,80,160,240,320\right\}$ for PubMed. The attack method used is DPGBA. The drop ratio $\beta=0.5$, and the number of iterations for random edge dropping are set as $K=10$ for Cora and $K=20$ for PubMed. The results are shown in Table 11. From the table, we observe: (i) Though there are fluctuations in the recall and precision in detecting poisoned nodes, the precision always exceeds $85\%$. This indicates that our method of random edge dropping consistently leads to higher prediction variance, helping us precisely identify poisoned nodes regardless of the number of triggers. (ii) With the different number of triggers, our method demonstrate consistently superior performance in defense, with ASR close to $0\%$ and ACC comparable to Clean ACC. Notably, when the number of triggers is $240$ for PubMed, the detection recall is around $66\%$, but we still achieve good results in terms of ASR and ACC. This further indicates the stability of our method in training a robust node classifier using Eq. 6, even if some poisoned nodes are not identified. Additional results on the visualization of prediction variance for poisoned nodes and clean nodes with different numbers of triggers are provided in Fig. 11 and Fig. 12.
@@
Appendix P Additional Results of Defense Performance and Ability to Detect Poison Nodes with Trigger Size as 1
In this section, we conduct experiments to investigate how different trigger sizes impact the performance of RIGBD in terms of backdoor defense and poisoned node detection. Specifically, we set the trigger size to 1 and conducted experiments on the Cora, Citeseer, Pubmed, and ogbn-arxiv datasets, keeping all other settings the same as in Section 5.2. The results are shown in Table 12, demonstrating that RIGBD achieves strong performance even with a trigger size of 1.
Table 12: Result of backdoor defense with trigger size as 1
Dataset Method ASR ACC Recall Precision
Cora UGBA 92.53 82.22 - -
RIGBD 0.44 83.70 83.9 92.9
Citeseer UGBA 95.24 74.69 - -
RIGBD 0.00 73.20 98.7 98.7
Pubmed UGBA 92.58 84.58 - -
RIGBD 0.32 84.83 94.0 97.5
OGB-arxiv UGBA 85.29 65.38 - -
RIGBD 0.00 65.92 92.7 99.8
Appendix Q Reproducibility
Experimental and implementation details are in Sec. 5.1 and Appendix H. We provide a detailed algorithm description in Algo. 1. The code for RIGBD is publicly available at: github.com/zzwjames/RIGBD.
Appendix R Limitations and Future Work
Our proposed framework aims to defend against graph backdoor attacks during the training phase. We believe it is also important to develop new techniques to defend against graph backdoor attacks during the inference stage, which we leave for future work. As the first work targeting all existing representative graph backdoor attacks, we hope to inspire future development of graph backdoor defense algorithms. Additionally, this paper focuses exclusively on graph-structured data, and it would be interesting to explore extensions to other domains. Given the nature of this work, there are no easily predictable negative social impacts.